# Scaffolding Creativity: Integrating Generative AI Tools and Real-world Experiences in Business Education

### v1

@@ Abstract.
This study explores the integration of Generative AI tools and real-world experiences in business education. Through a case study of an innovative undergraduate course, we investigate how AI-assisted learning, combined with experiential components, impacts students’ creative processes and learning outcomes. Our findings reveal that this integrated approach accelerates knowledge acquisition, enables students to overcome traditional creative barriers, and facilitates a dynamic interplay between AI-generated insights and real-world observations. The study also highlights challenges, including the need for instructors with high AI literacy and the rapid evolution of AI tools creating a moving target for curriculum design. These insights contribute to the growing body of literature on AI in education and provide actionable recommendations for educators preparing students for the complexities of modern business environments.
copyright: noneconference: ; ; copyright: none
@@ 1. INTRODUCTION
The integration of Artificial Intelligence (AI) in education has rapidly evolved, yet its application in creative and entrepreneurial education remains underexplored. While Generative AI tools have shown promise in assisting with tasks like writing and coding (Godwin-Jones, 2022; Alharbi, 2023; Becker et al., 2023), their role in fostering creativity and business acumen in educational contexts is still is in an early stage of active research and development. While current research explores diverse applications of AI in education, including but not limited to personalized learning systems, automated assessment tools, intelligent tutoring systems(Limo et al., 2023; Fuchs, 2023; Kim et al., 2019), little has been explored in the potential to scaffold complex, interdisciplinary learning experiences.
Simultaneously, business education faces significant challenges in keeping pace with the rapidly evolving landscape of AI and its implications for the business world. This lag is exacerbated by the scarcity of pedagogical resources tailored for business students(Xu and Babaian, 2021) , contributing to what has been described as ’curriculum chaos’ (Bunch, 2020). As Sollosy and McInerney (Sollosy and McInerney, 2022) note, changing curriculum is often a ’Herculean task,’ with the rapid pace of change in the business world leading to equally rapid and sometimes disorienting changes in management education.
The interdisciplinary nature of AI also often conflicts with the traditional, discipline-by-discipline approach often employed in business schools (Currie and Knights, 2003). This creates a disconnect between the siloed nature of business education and the integrative demands of AI in business practice, potentially leaving graduates ill-prepared for the dynamic challenges of an AI-influenced business environment.
@@
The challenge is further compounded by vast variations in students’ AI literacy (Hornberger et al., 2023) and their ability to effectively utilize rapidly evolving AI tools. Novice users, in particular, face significant hurdles in prompt designs (Zamfirescu-Pereira et al., 2023). The rapid rollout of advanced language models like GPT-4o, and Claude 3.5 Sonnet, along with image and video generation models such as Midjourney and Runway, presents a moving target for educators. While students often have some level of access to these tools (e.g., free daily messages), their proficiency in leveraging these technologies for creative and entrepreneurial tasks varies widely. Moreover, most current business educational research has focused primarily on text-based models, leaving a significant gap in understanding how to effectively integrate multi-modal AI tools into business education.
These deficiencies highlight the crucial role of instructors in guiding students’ interaction with AI. Educators must not only teach the technical aspects of AI tools but also foster the development of prerequisite skills such as critical thinking, digital literacy, and media literacy (Chiu et al., 2022). This complex educational landscape raises critical questions for instructors about the optimal level of scaffolding and how to strike the right balance between AI assistance and human creativity in the learning process. This also requires a shift towards more interdisciplinary teaching approaches (Chiu et al., 2022; Chiu, 2023) that reflect the integrative nature of AI in business.
To address these issues, this case study reports on the student learning processes and outcomes in an innovative undergraduate business course that combines AI-assisted learning with real-world experiences, through an interdisciplinary breadth-over-depth teaching approach.
We find that this integrated approach:
@@
    Accelerates knowledge acquisition and content creation, enabling students to produce high-quality outputs at an unprecedented pace.
@@
    Allows business students to overcome traditional creative skill barriers (such as design and copywriting) that previously limited their ability to produce comprehensive, multi-faceted business outputs beyond a PowerPoint Presentation.
@@
    Facilitates a dynamic interplay between AI-generated knowledge and insights and real-world observations and validation.
  4. (4)
    Complements the siloed and fragmented educational experiences.
  5. (5)
    Encourages greater expression of individual ideas in projects.
We also find further challenges:
  1. (1)
    There is now an increased demand for instructors with high AI literacy, interdisciplinary expertise, and relevant localized knowledge.
  2. (2)
    Rapidly evolving AI tools create a moving target for curriculum design, questioning the sustainability of current educational models.
  3. (3)
    Traditional pedagogical timelines and assessment methods are challenged by accelerated learning processes and breadth-over-depth approaches.
  4. (4)
    Logistical and scalability issues of immersive real-world components may limit widespread adoption.
Overall, our findings not only contribute to the growing body of literature on AI in education but also provide actionable insights for educators seeking to prepare students for the complexities of modern business environments. Furthermore, our insights challenges traditional pedagogical models, suggesting new directions for innovative interdisciplinary curriculum designs that fosters creativity and practical skills in the AI-augmented business landscape.
@@ 2.1. Course Overview
The course Reinventing the Brand was originally designed for undergraduate students to apply branding and user-centered design principles in reimagining existing well-known brands. However, recognizing the increasing prevalence of individual-created brands on social media and the decreasing costs of AI-assisted content creation, the course evolved to focus on students inventing their own brands. This shift also recognized the potential for AI to reshape the business landscape, potentially enabling small business owners to flourish. By having students create their own brands, the course aimed to prepare them for a future where AI might lower branding barriers to entry in various industries.
The curriculum allowed students to leverage their existing knowledge while exploring the practical challenges of brand creation in an AI-augmented environment. Students are required to produce tangible outputs including brand websites with logos and designs, social media posts, presentations detailing brand development rationales, AI-generated commercial videos, and physical artifacts such as packaging prototypes and fabric samples.
This curriculum design reflects a deliberate breadth-over-depth approach, aimed at providing students with a comprehensive overview of the multifaceted nature of brand creation and management in the digital age. The course structure is designed to break down the traditional discipline silos in business education and allow for personalized interdisciplinary learning trajectories.
2.2. Participants
@@
  • Note: S7 and S8 completed the final project as a group.
@@ 2.3. Phased Integration of AI Tools and Real-World Experiences
Students were encouraged to use ChatGPT throughout the course. At the beginning of the semester, most students used ChatGPT 3.5, as GPT-4o was not freely accessible until mid-semester (capped at 5 free messages per day for the free tier). The use of ChatGPT was optional, leading to varied adoption rates among students, with many choosing not to use it at the beginning.
Table 2. Case Study Focus
Scope of Focus Details Sample size (N)
Mini-Presentation 4 (MP4) Research on raw materials and production 8
MP4 ChatGPT documentation Prompt history from three students 3
Field Trip to Yiwu Observation and reflection session 8
Branding Proposal Initial proposal for the brand concepts (text-based)a 7
Midjourney documentation Interaction history from 4 Students 4
  • a
    Students were encouraged to submit their prompt histories but are not required, resulting in an even smaller sample size for ChatGPT and Midjourney prompt analyses.
@@
To facilitate the creation of tangible outputs such as brand images and AI video commercials, we introduced a Basic GenAI training session in Week 9. This session covered tools such as Midjourney, Leonardo.AI and Pika. For logo creation and voiceover, students utilized various third-party specialized AI-powered tools.
@@ 2.4. Scope of Focus and Research Methodology
This study employed a multi-phase, mixed-methods approach to investigate the integration of AI tools in business education. Despite the small sample size (N=8), the depth and richness of data collected across multiple phases allowed for a comprehensive exploration of AI-assisted learning in this context.
@@
Given the exploratory nature of the study and the diverse data collected, our analysis focused on three key areas (Table 2):
@@
    Integration of AI insights with real-world experiences: We compared students’ AI-generated knowledge before the Yiwu trip with their post-trip understanding. This analysis was based on observational notes from the field trip and discussions with students during and after the visit.
@@
Limitations of Small Sample Size
Although students were highly encouraged to submit their entire prompt histories, we received only a handful of submissions. While this small sample size limits our ability to draw broad conclusions from the prompt analysis, this study serves as a foundation for future, larger-scale investigations into the integration of AI tools in educational settings.
@@ 3. FINDINGS
Our analysis focuses on three components of the course: Mini-Presentation 4 (MP4), the Yiwu field trip, and students’ use of Midjourney for their final presentations and AI video commercials. These components provide many insights into how students leverage text-based AI tools for research, integrate real-world experiences, and use multimodal tools as creative scaffolds for visual transformation.
Table 3. Student Prompts and Observed Behaviors

### v2

@@ Abstract.
This exploratory study investigates the intersection of Generative AI tools and experiential learning in business education. Through a case study of an innovative undergraduate course, we examine how students interact with and adapt to various AI modalities—from text-based tools to image generation—alongside real-world experiences. Our findings reveal how this integrated approach enables novice users to overcome creative barriers, accelerates skill acquisition, and creates a dynamic interplay between AI-generated insights and real-world validation. We identify critical interaction challenges, including prompt engineering patterns and the need for more intuitive AI interfaces in educational contexts. These insights inform the design of future AI tools for creative learning and contribute to broader HCI discussions about human-AI collaboration in educational settings.
AI-assisted learning, Artificial Intelligence in Education (AIEd), Higher Education, Experiential learning, ChatGPT, Midjourney
This is the author’s version of the work accepted for publication in CHI EA ’25: Extended Abstracts of the CHI Conference on Human Factors in Computing Systems. The final published version is available at https://doi.org/10.1145/3706599.3720283
conference:  ;  ;  ccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
The rapid evolution of Generative AI tools is fundamentally transforming how we approach teaching and learning across disciplines. While current research has explored various applications of AI in education, including personalized learning systems and automated assessment tools(Limo et al., 2023; Fuchs, 2023; Mizumoto and Eguchi, 2023), little attention has been paid to how these tools can scaffold complex, interdisciplinary learning experiences. This gap is particularly evident in business education, where students must increasingly navigate the intersection of creativity, technology, and business strategy.
The challenge is multifaceted: students need to master not only business fundamentals but also develop the creative and technical capabilities demanded by an AI-augmented business environment. Traditional business education, with its discipline-by-discipline approach(Currie and Knights, 2003), creates a disconnect between siloed learning and the integrative nature of real-world practice. This disconnect is further amplified by vast variations in students’ AI literacy(Hornberger et al., 2023) and their ability to effectively leverage rapidly evolving AI tools in creative and strategic tasks.
@@
Novice users also face significant hurdles in prompt engineering (Zamfirescu-Pereira et al., 2023; Šedlbauer et al., 2024). The rapid deployment of advanced language models like GPT-4 and Claude 3.5 Sonnet, alongside image and video generation models such as Midjourney and Runway, presents educators with a moving target. While students often have access to these tools, their proficiency in leveraging them for creative and entrepreneurial tasks varies widely. Moreover, most educational research has focused primarily on text-based models, leaving a significant gap in understanding how to effectively integrate multimodal AI tools into experiential learning.
To address these challenges, this case study serves as exploratory research to understand how the integration of AI tools with real-world experiences impacts students’ creative processes and learning outcomes in business education. While the study does not employ formal pre- and post-assessments or control groups due to the nature of the course setting, and the sample is limited, this initial investigation aims to identify promising directions for future research in this rapidly evolving space:
@@
    How students leverage different AI tools to overcome creative barriers
@@
    The interplay between AI-generated insights and real-world observations
@@
    The role of AI tools in scaffolding the development of visual and creative skills
Creativity, broadly defined as the ability to generate ideas that are both novel and effective (Runco and Jaeger, 2012), plays a central role in this study, as AI tools have the potential to redefine how students engage with idea generation, problem-solving, and execution. Business creativity is often market-oriented, manifesting through strategies, business models, and market solutions (Berglund and Wennberg, 2006). This historically makes it more challenging to evaluate creativity’s effectiveness in real-world contexts. In this study, we redefine ‘effectiveness’ in creativity not only as practical feasibility but also as the ability to bridge conceptual thinking with real-world execution, with the aid of AI tools.
Our findings reveal that this integrated approach not only accelerates knowledge acquisition but also enables students to overcome traditional creative barriers, create more tangible non-text-based deliverables, and develop more nuanced understanding of their business ideas through the dynamic interaction between AI tools and real-world experiences. We also identify key challenges, including the need for new assessment methods and the rapid evolution of AI capabilities creating a moving target for curriculum design.
These insights contribute to ongoing HCI discussions about human-AI collaboration in learning contexts and provide actionable recommendations for designing educational experiences that effectively combine AI tools with real-world practice. As an exploratory study, our work not only documents initial findings in this emerging space but also aims to catalyze future research directions in the design of AI interfaces that better support creative learning and interdisciplinary skill development. The patterns and challenges identified here, while preliminary, point to rich opportunities for more extensive investigation by the HCI community.
@@ 2.1. Course Overview
The course Reinventing the Brand evolved from initially focusing on reimagining existing brands to having students create their own brands, recognizing the increasing prevalence of individual-created brands and the potential of AI to lower barriers to entry. This shift aimed to prepare students for a future where AI might enable small business owners to compete more effectively in branding and content creation (West et al., 2018; Wang et al., 2023).
Students were required to produce tangible outputs including:brand websites with logos and designs, social media posts and content strategies, brand development presentations, AI-generated commercial videos and physical artifacts (eg. packaging prototypes, fabric samples).
The curriculum adopted a breadth-over-depth approach to provide a comprehensive overview of brand creation and management in the AI era. This design deliberately broke down traditional business education silos, allowing for personalized interdisciplinary learning trajectories that reflect the integrative nature of modern brand development.
@@
*S7 and S8 completed the final project as a group.
2.2. Participants
@@ 2.3. Phased Integration of AI Tools and Real-World Experiences
Students were encouraged to use ChatGPT throughout the course. At the beginning of the semester, most students used ChatGPT 3.5, as GPT-4o was not freely accessible until mid-semester (capped at 5 free messages per day for the free tier). The use of ChatGPT was optional, leading to varied adoption rates among the students.
@@
To facilitate the creation of tangible outputs such as brand images and AI video commercials, we introduced a Basic GenAI training session in Week 9. This session covered coomercial AI tools such as Midjourney, Leonardo.AI, Pika, Suno.ai, ElevenLabs.
Table 2. Case Study Focus
Scope of Focus Details Sample size (N)
Mini-Presentation 4 (MP4) Research on raw materials and production 8
MP4 ChatGPT documentation Prompt history from three students 3
Field Trip to Yiwu Observation and reflection session 8
Branding Proposal Initial proposal for the brand concepts (text-based)111a 7
Midjourney documentation Interaction history from 4 Students 4
1 Students were encouraged to submit their prompt histories but are not required, resulting in an even smaller sample size for ChatGPT and Midjourney prompt analyses.
@@ 2.4. Scope of Focus and Research Methodology
Given the exploratory nature of this study, we documented student interactions with AI tools across the following components (Table 2):
@@
    Integration of AI insights with real-world experiences: We compared students’ AI-generated knowledge before the Yiwu trip with their post-trip in-class reflections, based on observational notes and discussions with students before, during, and after the visit.
@@
Data Analysis Approach and Limitations
Due to the constraints of conducting research within an active course setting, formal experimental designs such as pre/post assessments or control groups were not feasible. Additionally, given the small sample size and the exploratory nature of this study, we did not employ structured qualitative analysis methods such as thematic analysis or inductive coding. Instead, our findings were drawn from observed patterns, student-reported experiences, and qualitative comparisons of AI-assisted outputs and non-AI-assisted outputs in previous similar courses. Specifically:
  • Prompt Histories were analyzed through instructor observations of how students refined their prompts and iterated on AI-generated content. Given the small sample size, this analysis remains an interpretative assessment rather than a formally coded dataset. Some students explicitly described their decision-making process in reflections, which provided additional qualitative insights.
  • Project Output were reviewed to identify differences in creativity and execution. While this study does not conduct a formal comparative analysis, the instructor qualitatively compared branding portfolios from the current AI-assisted course with branding portfolios from a prior similar course that did not use AI. However, these insights are based on instructor experience rather than controlled experimental comparisons, and future studies could incorporate structured comparative methods to assess AI’s impact on creative capacity more systematically.
  • Classroom Observations & Discussions were informally documented through instructor interactions with students during the course. However, these insights were not systematically logged, as they emerged through ongoing discussions rather than structured data collection. Future research will aim to capture these observations more systematically through structured field notes or follow-up interviews.
While this study does not formally measure whether AI enhanced students’ creative capacity or merely improved execution efficiency, instructor observations suggest that students engaged AI for both ideation and execution. However, the extent and impact of these effects remain unclear and require further research.
@@ 3. FINDINGS
3.1. AI-Assisted Research in Unfamiliar Topics
The assignment instructions included general ChatGPT usage principles (Appendix B):
  • encouraged open-ended exploration
  • start with broad questions to establish a foundation
  • take an iterative approach
While all students used ChatGPT-3.5 for research, only three documented their prompt histories, providing some insights into different learning patterns:
Non-Linear Learning Trajectories and Knowledge Integration
Despite the assignment suggesting a structured approach from broad to specific inquiries (Appendix B), students demonstrated distinctly different patterns of AI interaction based on their domain knowledge. These patterns reveal how students adaptively integrated AI assistance with their existing expertise and external resources.
Students with domain expertise demonstrated sophisticated, non-linear approaches to knowledge building. For instance, S1, leveraging garment production experience, began with specific material inquiries (”Which items of clothing can be made with 100% hemp fabric?”[Appendix A, S1, P1]) rather than broad material exploration. This approach indicates an attempt to validate and expand existing knowledge using ChatGPT as a reference point.
In contrast, S2 drawing on design experience, initiated the interaction by describing her product vision with clear design-oriented specifications such as “water bottle pockets” and “enough space for students”[S2,P1]. When this prompt didn’t yield the desired response, S2 adapted by returning to broader material questions.
S3, who had limited domain experience, followed a more systematic approach aligned with the assignment suggestions. This pattern suggests that students without prior domain knowledge may rely more heavily on the structured approach provided in the instructions.
Students’ background in domain knowledge also seemingly affected their analytical approach when interacting with ChatGPT. Students with less domain knowledge tend to accept AI responses more readily, using them as scaffolding to build general understanding. But the students with certain domain knowledge are more likely to demonstrate a more critical and nuanced engagement with AI-generated information. For example, when S1 learned that hemp can be used to create “even jeans” she immediately followed up with a question [S1,P4] about the differences between cotton and hemp jeans for deeper inquiry. Similarly, S2 cited information from internet sources[S2,P4], particularly when ChatGPT’s answers seemed too theoretical for her practical, production-oriented questions.
Finally, S1 demonstrated a tendency to inquire about tangential knowledge. In final queries, S1 focused on ramie [S1, P7-P8], a fabric not mentioned in previous responses. This sudden shift in focus could represent either a recollection of prior knowledge or genuine curiosity sparked by the AI interaction. This behavior could suggest the AI’s responses might trigger memories of related concepts or introduce new ideas that pique the student’s interest, leading to unexpected learning paths.
The diverse learning trajectories observed in these cases illustrate how AI tools can facilitate non-linear, adaptive learning processes by enabling students to dynamically integrate new information with prior knowledge and external resources (Rasul et al., 2023; Rudolph et al., 2023). These interactions reveal the potential of AI to scaffold both novice and expert learners. However, the same flexibility that supports adaptive learning also introduces the risk of scattered focus if learners lack guidance. This highlights the importance of designing AI systems that not only provide information but also guide users in structuring their inquiry and critically evaluating responses.
Accelerated Domain-Specific Knowledge Acquisition
This observation demonstrates the potential of AI as a tool for just-in-time learning in project-based contexts (Seo et al., 2021; Kim and Adlof, 2024), where students can quickly acquire specific expertise needed for their individual projects within compressed timelines. However, this accelerated knowledge acquisition raises important considerations about the balance between speed, depth and retention in learning (Rudolph et al., 2023; Seo et al., 2021).
3.2. Dynamic Interaction with Real-World Experiences
Our integration of AI tools with real-world experiences revealed interesting patterns in how students navigate between digital and physical learning environments, particularly in validating and adapting AI-generated knowledge.
Human-AI Collaboration in Complex Environments
After their AI-assisted research into raw materials and production processes, students successfully identified potential manufacturers on 1688.com, a major B2B e-commerce platform, demonstrating how AI can expand novice users’ capabilities in complex market research tasks typically reserved for experienced professionals.
A significant insight emerged from how students combined AI-generated knowledge with real-world experiences in the Yiwu market. The rapid knowledge acquisition facilitated by AI enabled students to effectively engage in vendor interactions, with some successfully presenting themselves as experienced buyers. This was particularly notable as vendors typically identify and offer unfavorable deals to novice buyers who don’t ask the ”right” questions. The AI-assisted preparation seemed to have provided students with sufficient domain knowledge to ask pertinent questions and engage confidently in nuanced conversations.
The interplay between AI preparation and human adaptability suggests a novel approach in business experiential learning. Students used AI-generated knowledge as a foundation, which they dynamically adjusted based on real-time observations and interactions. This continuous recalibration between AI outputs and human experiences reflects ongoing discussions on Human-AI collaboration in learning contexts (Nah et al., 2023).
Uncovering AI limitations in Real-World Engagement
However, the real-world engagement also revealed limitations in AI-generated knowledge. For example, S1’s discovery that 100% hemp production was far less common than AI had implied. Moreover, many students encountered a wide array of materials and production techniques during their visit that were not mentioned in their AI-assisted research.
This discovery points to two known limitations of AI-generated knowledge: first, its tendency to provide generic answers unless prompted with very specific context (Memarian and Doleck, 2023), which is difficult for student users; and second, its potential lag in incorporating the latest innovations or highly specialized industry knowledge (Zhou et al., 2023). This suggests a need for AI interfaces that can incorporate real-world validation mechanisms and support iterative refinement of knowledge based on user feedback.
3.3. AI-Assisted Visual Transformation
Traditionally, business school assignments have relied on text-based presentations supplemented with basic visuals, partly due to students’ non-design backgrounds. Our study investigates how the introduction of multimodal AI tools transforms students’ ability to visualize and communicate business concepts. Four students (N=4) maintained detailed records of their Midjourney interactions in the class Discord server after the training session, while others used various AI image generation tools through personal accounts. Table 3 presents the statistics of these Midjourney interactions, including initial concept prompts, iterative prompts (usually style or component changes), upscaling actions, remixes, uploads, image-to-text descriptions and command error corrections.
AI as a Cognitive Scaffold and Human-AI Co-Creation
Our study revealed that AI tools functioned not merely as implementation aids, but as collaborative partners in both concept development and visualization. This partnership manifested as an iterative process where AI-generated visuals influenced conceptual thinking, which in turn guided further visual exploration. In particular, the introduction of multimodal AI tools after midterm appeared to serve as cognitive scaffolds which was particularly evident in the expansion and refinement of initial ideas:
Similarly, S2’s interaction with AI revealed how unexpected elements in generated images could spark new creative directions. After noticing a blurry figure in one AI-generated image, she incorporated human figures in subsequent prompts, better representing her brand’s emphasis on ease of use.
This collaborative process aligns with the vision that AI tools act not just as implementers of human ideas, but as active collaborators in the creative process (Kim et al., 2022; Yin et al., 2023). It also resonates with the notion of AI as a coach or a colleague (Lubart, 2005), stimulating new directions of thought and encouraging students to reconsider and enhance their initial concepts.
Table 3. Midjourney Interactions (Post-Training)
ID S2 S3 S4 S5
Interactions 90 31 40 12
Concepts 14 10 12 4
Iterations 45 10 19 6
Effective Interactions per Concept 2.6 1.0 1.6 1.5
1 Students could make errors in their prompts due to their unfamiliarity with the command format in Midjourney’s server on Discord, causing ineffective interactions.
Democratization of Visual Design Skills
The rapid adaptation to visual communication tools suggests that multimodal AI is democratizing design capabilities in non-design disciplines(Adetayo, 2024; Yin et al., 2023). All students, regardless of prior design experience, demonstrated strong capability in visual communication, producing professional-quality outputs that markedly exceeded typical business course submissions.
This democratization was particularly notable given that none of the students had prior experience with text-to-image or image-to-video tools in professional contexts before the Week 9 training session. The rapid skill acquisition suggests that current AI interfaces may be tapping into an innate capacity for visual expression previously constrained by technical barriers. This raises important implications for interface design. How might AI tools better support the transition from textual to visual thinking? What interface elements best facilitate creative exploration for non-designers? How can AI tools maintain professional output quality while supporting user agency and learning?
Shift Towards Individual Expression and Engagement
Despite encouragement to form pairs after the brand proposal stage, six out of eight students insisted on working individually. The ability to rapidly generate and refine designs independently may have reduced the perceived necessity of collaboration, as students no longer relied on peers to fill skill gaps. This aligns with the concept of AI as an Empowering Partner(UNESCO, 2023), in which AI tools act as individualized creative supports, fostering a sense of agency and competence. Additionally, this high motivation may resonate with Csikszentmihalyi’s Flow Theory(Csikszentmihalyi and LeFevre, 1989), suggesting that AI tools can potentially create an optimal challenge-skill balance that facilitates a state of flow where students felt both capable and motivated to pursue ambitious individual projects.
However, this shift toward individual learning raises important considerations for business education, where teamwork and interpersonal skills are crucial for professional success (Rasul et al., 2023). While AI enhances technical execution and creative iteration, it does not inherently foster the interpersonal abilities necessary for collaboration, negotiation, and shared decision-making—skills that traditionally develop through direct peer interaction. This suggests that AI-driven workflows may unintentionally deprioritize teamwork, potentially altering how students experience and develop collaborative competencies. Future course designs should explore how AI can be integrated without diminishing opportunities for interpersonal learning and teamwork.
4. IMPLICATIONS AND LIMITATIONS
The emergence of non-linear learning patterns in AI interactions suggests the need for more adaptive interfaces that can accommodate diverse learning approaches. Our observations reveal that users navigate AI capabilities in more complex ways, with their domain knowledge and creative goals significantly influencing their interaction patterns. This finding calls for new interface paradigms that can dynamically adjust to users’ varying expertise levels and learning trajectories. Such adaptive interfaces are particularly relevant for the design of personalized (or self-directed) learning systems, which currently tend to focus primarily on AI model capabilities and knowledge integration rather than supporting the complex, non-linear nature of creative exploration and skill development. Future AI systems could proactively suggest real-world tasks or experiments tailored to a user’s specific progress and interactions.
The interplay between AI-generated knowledge and real-world validation experiences suggests an important direction for future interface development. Current AI systems largely operate as closed loops, with limited ability to incorporate real-world feedback. Our findings indicate a need for interfaces that can not only integrate observational data from real-world experiences but also adapt iteratively based on this feedback. This iterative refinement is especially critical in educational contexts where theoretical knowledge must align with practical realities. Moreover, the adaptive and iterative interaction patterns observed in this study could inform the design of AI tools for professional and lifelong learning, where learners must continuously apply knowledge in evolving real-world scenarios.
From a broader perspective, our findings can be situated within Wu et al.’s Human-AI Co-Creation Model(Wu et al., 2021). It provides a useful conceptual foundation for understanding how humans and AI interact through cycles of perception, thinking, expression, collaboration, building, and testing. For instance, AI facilitates deeper and wider exploration during the thinking and expression phases, offering adaptive learning that supports ideation and concept development. This is then followed by real-world validation, where students build and test their knowledge by integrating AI-generated insights into practical observations and applications. Currently, AI’s role in perception remains limited, as it lacks the sensorimotor capabilities to directly interpret real-world contexts. Furthermore, while the authors of the model argue that humans and AI complement each other in the co-creative process, their relative strengths may shift across different learning contexts and evolve over time, Recognizing these complexities can inform the design of future AI learning systems that better support adaptive, iterative, and co-creative learning experiences.
Another promising area for future research emerges from our observations of individual creative work. The unexpected preference for individual work, despite the project’s complexity, suggests that AI tools may be reshaping the role of collaboration in learning environments. While AI-driven workflows offer significant advantages, they may also unintentionally limit opportunities for students to develop interpersonal skills essential for professional collaboration.Future research should explore hybrid approaches to ensure that students benefit from both personalized AI assistance and human-centered collaborative learning experiences.
The methodological limitations of this study also point to opportunities for further investigation. Due to the constraints of conducting research within an active course setting, this study is exploratory in nature and does not employ formal qualitative coding techniques or controlled comparative analysis. Future research should incorporate structured analysis and even longitudinal assessment framework to examine AI’s impact on conceptual learning, critical thinking, and business knowledge development more systematically.
Given the small sample size and single-course context, further research should also examine how human-AI creative collaboration unfolds across different educational settings and disciplines. Additionally, the observed challenges with prompt engineering and AI bias—particularly in image generation—underscore the need for more sophisticated interface mechanisms that help users better articulate their creative intentions and critically evaluate AI outputs.
Finally, the study did not assess how mismatches between AI-generated outputs and students’ expectations impacted their trust in AI-generated knowledge. Future research should investigate how students develop AI literacy skills, including their ability to assess AI reliability, recognize biases, and refine AI-generated content critically (Lee, 2024; Kim and Adlof, 2024).
These findings contribute to an emerging research agenda for the HCI community focused on developing more nuanced and effective approaches to human-AI collaboration in creative learning contexts. As educational institutions continue to embrace the integration of rapidly evolving AI tools, the development of interfaces that can effectively support these complex learning processes becomes increasingly critical.
Appendix A Student Prompt Histories and Observed Behaviors